OpenStudy (anonymous):

how do i find period: y=sin(5x +11\pi) -4

OpenStudy (campbell_st):

the coefficient of x is used to find the period given y = asin(bx) a= amplitude and the period is \[period = \frac{2\pi}{b}\] the phase shift is \[\frac{11}{\pi}\] and the whole graph is moved down to centre on y = -4 hope it helps

OpenStudy (anonymous):

thanks